Show that if n | m, where n and m are integers greater than 1, and if a≡b (mod m), where a and b are integers, then a≡b (mod n).




1
Expert's answer
2021-10-06T14:47:08-0400

If "n|m" , then there is integer number "k" such that "m=kn" .

The following conditions are equivalent: "a\equiv b\mod m" and "m|(a-b)" .

If "m|(a-b)" , then there is integer number "l" such that "a-b=ml" .

So, "a-b=ml=kn\cdot l= (kl)\cdot n" .

It means, that "n|(a-b)" or it can be written as "a\equiv b\mod n"